OpenStudy (anonymous):

PLEASE HELP ME Mega Movies hosted a film premiere on Friday night. They charged $8 for adults and $5 for children. One hundred thirty adults and children attended, and $986 was made in ticket sales. How many children and how many adults went to the film premiere? 15 children; 115 adults 18 children; 112 adults 12 children; 118 adults 22 children; 108 adults

OpenStudy (anonymous):

a - adults c - children we have: 8*a + 5*c = 986 a+c=130 so we can solve for c: a= 130-c 8*(130-c) + 5*c = 986 1040-8c+5c=986 1040-3c=986 c=18 so b.

OpenStudy (anonymous):

thanks